OpenStudy (abbles):

How to multiply (x - (3 + 2i)) times (x - (x - 2i)) ?

OpenStudy (abbles):

\[(x - (3 + 2i))(x - (3 - 2i))\]

OpenStudy (mathmate):

You can start with FOIL \((a-B)(a-b)=a^2-aB-ab-Bb\) a=x, B=3+2i, b=3-2i then you distribute aB as x(3+2i), and ab as x(3-2i). Finally, \(bB = (3+2i)(3-2i)\) can be expanded to the difference of two squares, namely \((3+2i)(3-2i)=3^2-(2i)^2.\) Recall: \((2i)^2=2^2i^2\)

OpenStudy (mathmate):

* \((a-B)(a-b)=a^2-aB-ab\color{red}{+}Bb\)

OpenStudy (abbles):

Thanks - very helpful. I'll go through a few similar problems and see if I hit any walls.